OpenStudy (mef4.0):

simplify the expression cos^2 A/1 + sin A

myininaya (myininaya):

\[\cos^2(A)=1-\sin^2(A)=(1+\sin(A))(1-\sin(A))\]

myininaya (myininaya):

that should help

OpenStudy (lalaly):

\[\large{\cos^2A=1-\sin^2A}\]\[\large{1-\sin^2A}=(1-sinA)(1+sinA)\] \[\large{\frac{\cos^2A}{1+sinA}}=\frac{1-\sin^2A}{1+sinA}= \frac{(1-sinA)\cancel{(1+sinA)}}{\cancel{1+sinA}}\]

OpenStudy (anonymous):

= 1 - sin^2A / 1 + sinA = (1-sinA)(1+sinA) / (1+ sin A) - nearly done

OpenStudy (mef4.0):

Thank you all